# Thermodynamic Signatures of Gaussian Entanglement Beyond Entropy

### v1

@@ Abstract
Figure 1: Difference between $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ and $\Delta\mathcal{E}_{\text{rel}}$ for the parametric family of Gaussian two-mode squeezed states, as a function of the squeezing parameter $z$ and the mean fluctuations factor $k$. Values of fluctuation difference $\gamma$ have been set to {1,0.5,0} (left to right). The dashed and solid red lines represent the boundaries between separable and entangled states according to our criterion and the PPT condition, respectively. As predicted by Corollary 1, these boundaries coincide when $\gamma=0$.
Non-Gaussian states.– Gaussian states are experimentally and theoretically important, but their ease of characterization belies fundamental limitations in their practical usefulness for some quantum information tasks [51]. This is what motivates our interest in exploring the non-Gaussian realm [35] and to answer the question of how far the relationship between the REG and (quantum) correlations can be extended into it. By default, the results above (Theorems 1 and 2) directly apply to the class of non-Gaussian states that can be produced through local unitary operations on Gaussian ones. Outside this class, we examine the example of one photon-subtracted TMS states, whose REG can be computed from Eqs.3,5, and 6, exploiting the form of the covariance matrix derived in [52] (detailed derivations can be found in Appendix D). In Figure 2, $\Delta\mathcal{E}_{\text{rel}}$ is represented as a function of the same squeezing and fluctuation parameters $z,k$, respectively, of the initial TMS Gaussian state prior to the photon subtraction. Regarding the separability condition (dashed line in the figure), in the non-Gaussian case we use the criterion given by Shchukin and Vogel [36] as a benchmark, whose negativity is sufficient for entanglement. In particular, from the infinite set of inequalities given by such criterion, we pick the following:
$\langle\hat{a}^{\dagger}\hat{a}\rangle_{\rho}\langle\hat{b}^{\dagger}\hat{b}\rangle_{\rho}-\langle\hat{a}\hat{b}\rangle_{\rho}\langle\hat{a}^{\dagger}\hat{b}^{\dagger}\rangle_{\rho}<0\Longrightarrow\ \rho\ \text{entangled}$ (10)
where $\hat{a},\hat{b}$ are the annihilation operators associated to modes $A$ and $B$, respectively.
We observe numerically that there exists some trend by which entangled states (on the left of the separability boundary) display higher values of $\Delta\mathcal{E}_{\text{rel}}$, which increase the further they lie from the criterion’s dashed line. This behavior could allow us to identify the highest value of $\Delta\mathcal{E}_{\text{rel}}$ within the separability regime and establish it as an entanglement witness. We thus conjecture that this link between REG and entanglement observed in TMS photon-subtracted states could potentially be extended to other different classes of non-Gaussian states. The verification of this claim is left for future work.
@@
Figure 2: REG of photon-subtracted TMS($z,k$) as a function of the fluctuation parameter $k$ of the initial thermal state and the parameter $z$ of the Gaussian squeezing transformation. The dashed line indicates the region of entangled states witnessed by the Shchukin-Vogel (SV) condition.
@@
Discussions.– In this work we characterize correlations in CV systems via fundamental thermodynamic processes. Our results show that for bipartite Gaussian states, any correlation, even between asymmetric modes, ensures a strictly positive ergotropic gap, revealing a distinct form of correlation independent of entropy or informational quantities. Physically, this means that energy can be locked into correlations, making it inaccessible via local operations alone. For pure states, this phenomenon yields a valid entanglement quantifier, equivalent to the entropy of entanglement. However, for mixed states, the ergotropic gap diverges due to the unbounded energy spectrum of CV systems. To resolve this, we introduce the relative ergotropic gap (REG), which eliminates the divergence while preserving its role as a correlation indicator. This shows that entanglement shapes information and energy processing in a very different way for discrete and continuous systems. Notably, REG is functionally independent of mutual information, uncovering a distinct yet operationally meaningful aspect of quantum correlations. Furthermore, it enables a sharper classification of states, distinguishing entangled from separable ones through analytically derived bounds. While REG is not a universal entanglement measure, it provides a necessary and sufficient criterion under symmetric thermal conditions, paralleling the PPT test. When energy measurements are available [53], this could provide a more experimentally accessible entanglement criterion than full state tomography. Moreover, our results suggest REG’s broader applicability outside the Gaussian regime — an open direction for future study. Overall, our work establishes REG as an energy-based, experimentally accessible witness of quantum correlations, bridging thermodynamics and information-theoretic perspectives, and offering a novel tool for entanglement detection and characterization in near-term quantum optical technologies.
@@ Acknowledgements.
Acknowledgements: This project has been funded by the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program). F.C. acknowledges funding from the European Union (EQC, 101149233). M.A. acknowledges the funding supported by the European Union (Project- QURES, Grant Agreement No. 101153001). G.A. acknowledges funding from the UK Research and Innovation (UKRI EPSRC Grant No. EP/X010929/1). The views and opinions expressed in this Letter are, however, the sole responsibility of the author(s).
I Appendix A. Separability of Gaussian two-mode states
@@
$\sigma_{sf}=\begin{pmatrix}a&0&c_{1}&0\\ 0&a&0&c_{2}\\ c_{1}&0&b&0\\ 0&c_{2}&0&b\end{pmatrix},\quad\sigma_{A}=a\mathbb{1}_{2},\quad\sigma_{B}=b\mathbb{1}_{2},\quad\sigma_{AB}=\begin{pmatrix}c_{1}&0\\ 0&c_{2}\end{pmatrix},$
@@
II Appendix B. Proof of Theorem 1
@@
III Appendix C. Parametrization of Gaussian two-mode mixed states
@@
In what follows, we will adopt the $xpxp$ convention for mode ordering for the definition of the matrices appearing in the previous expression; that is, we will take the quadrature vector as $\hat{\vec{r}}=(\hat{x}_{A},\hat{p}_{A},\hat{x}_{B},\hat{p}_{B})^{T}$ as opposed to the alternative ordering $\hat{\vec{r}}=(\hat{x}_{A}\hat{x}_{B},\hat{p}_{A},\hat{p}_{B})^{T}$. The matrix
@@
$\omega\equiv\omega_{A}$: the natural oscillating frequency of mode $A$
@@
IV APPENDIX D. Calculation of relative ergotropic gaps
@@
IV.1 Gaussian two-mode states
@@
$\sigma_{\text{gp}}=G\ \sigma G^{T}$ (26)
@@
Substituting expressions 31,32 and 21 into Eq.30, we derive the explicit covariance matrix of the local passive state, whose energy, computed through Eq.3, yields:
E_lp = ω2 [ (k+γ)^2 cos^4θ+(k-γ)^2 sin^4θ+(k^2-γ^2)cos^2θsin^2θ(z12+ z22z1z2) + α(k-γ)^2 cos^4θ+(k+γ)^2 sin^4θ+(k^2-γ^2)cos^2θsin^2θ(z12+ z22z1z2) - (1+α) ]
@@
IV.2 Non-Gaussian example: photon subtracted TMS state
The computation of the Gaussian relative ergotropic gap for a non-Gaussian state is exactly analogous to that of a Gaussian state, and again requires knowledge of the state’s covariance matrix exclusively. Such CM is not accessible in general, but it is known for some specific classes of non-Gaussian states. As an example, we focus on two-mode squeezed (TMS) states that undergo a photon subtraction (non-Gaussian operation) on their first mode. The covariance matrix $\sigma$ of the resulting state $\rho$ is derived in [52]:
$\sigma=\sigma_{0}+2\frac{(\sigma_{0}-\mathds{1})P(\sigma_{0}-\mathds{1})}{\text{Tr}[(\sigma_{0}-\mathds{1})P]}$ (35)
$P=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$
The resulting $\sigma$ is in standard form (see Eq.4), with:
and its symplectic eigenvalues can be computed through
Finally, the expression for the relative ergotropic gap is directly derived from the above values and Eqs.5 and 6.
V APPENDIX E. Proof of Theorem 2 and Corollary 1
@@ IV.2 Non-Gaussian example: photon subtracted TMS state
Part (a). First we note that by expansion of the binomial squares $(k+\gamma)^{2}$ and $(k-\gamma)^{2}$, the sum of square roots in the expression of the relative ergotropic gap (Eq.34) yields the following structure
@@
The following figure illustrates the meaning of these two bounds. Relative ergotropic gap is shown for $N=5\times 10^{2}$ randomly generated two mode Gaussian states. The blue dashed line represents the minimum $\Delta\mathcal{E}_{\text{rel}}$ for entangled states, while the upper bound for separable states is displayed in dashed red. Since the two bounds $\mathcal{B}^{\text{ent}}_{\mathrm{min}}$ and $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ are dependent on $k,\gamma$ and $\alpha$, the values of these parameters have been fixed so a single threshold holds for all of the simulated states. Neither entanglement nor separability of states with REG lying in between the two lines can be witnessed by means of our criterion. However, any value of REG above the red line or below the blue one can unambiguously certify the state is entangled or separable, respectively.
@@
If $\rho_{G}$ has equal fluctuation factors $k_{A},k_{B}$ for both modes, then $k=k_{A}=k_{B}$ and $\gamma=0$. In that case, $\mathcal{B}^{\text{sep}}_{\mathrm{max}}=\mathcal{B}^{\text{ent}}_{\mathrm{min}}$. From part (a) of Theorem 2 we derive that Eq.9 is necessary for separability. Sufficiency follows from part (b). ∎
VI APPENDIX F. Independency of relative ergotropic gap

### v3

@@ Abstract
Figure 1: Difference between $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ and $\Delta\mathcal{E}_{\text{rel}}$ for the parametric family of Gaussian two-mode squeezed states, as a function of the squeezing parameter $z$ and the mean fluctuations factor $k$. Values of the fluctuation difference $\gamma$ have been set to {1, 0.5, 0} (left to right), and the frequency ratio is set at $\alpha=1$ for all cases. The dashed and solid red lines represent the boundaries between separable and entangled states according to our criterion [Eq. (7)] and the PPT condition, respectively. As predicted by Corollary 1, these boundaries coincide when $\gamma=0$.
Non-Gaussian states.– Gaussian states are both experimentally and theoretically important. However, their ease of characterization also comes with fundamental limitations regarding their usefulness in quantum information tasks [58]. This motivates our exploration of the non-Gaussian regime [41] and the question of how far the relationship between the ergotropic gap and quantum correlations can be extended. For generic non-Gaussian states, evaluating the ergotropic gap becomes highly challenging and may even be intractable. By contrast, the Gaussian ergotropic gap $\Delta\mathcal{E}_{G}$ [Eq. (2)] — together with the associated Gaussian relative ergotropic gap (G-REG) — remains fully determined by the covariance matrix, thereby retaining computational accessibility. This observation suggests that while $\Delta\mathcal{E}$ provides the most general characterization, the Gaussian counterpart $\Delta\mathcal{E}_{G}$ may serve as a practical proxy in situations where the exact computation is prohibitive. In what follows, we investigate how $\Delta\mathcal{E}$ (where calculable), $\Delta\mathcal{E}_{G}$, and their relative forms can act as witnesses of entanglement in different families of non-Gaussian states.
By construction, the general results derived earlier (Theorems 1 and 2) apply directly to non-Gaussian states obtained from Gaussian states through local unitary operations. However, the separable bound established in Eq. (7) cannot be straightforwardly generalized to arbitrary non-Gaussian states. To illustrate these subtleties, we now examine a specific class of states where $\Delta\mathcal{E}_{G}$ becomes trivial and fails as an entanglement witness, while $\Delta\mathcal{E}$ retains its detecting power.
Lemma 3.
Any symmetric superposition of bipartite Fock states, i.e., any state of the form $\frac{1}{\sqrt{2}}(|nm\rangle+|mn\rangle)$ with $n,m\in\mathbb{N}$, is passive under Gaussian operations.
The proof, provided in Appendix G.1 [49], shows that the covariance matrix of such states coincides with that of a Gaussian-passive state. Consequently, $\Delta\mathcal{E}_{G}=0$ even though the state is pure and entangled. Notably, however, the entanglement in this class can still be faithfully witnessed through $\Delta\mathcal{E}$, when ergotropy is defined with respect to arbitrary unitary operations.
[…]
Importantly, for the above class of Fock states, $\Delta\mathcal{E}$ can be easily evaluated. In contrast, for genuinely infinite-dimensional non-Gaussian states, one must rely on $\Delta\mathcal{E}_{G}$ to circumvent both experimental and theoretical complications. This raises the question whether nontrivial separability bounds for $\Delta\mathcal{E}_{G}$ can be established in such cases. Below, we provide an affirmative answer by considering an experimentally relevant family: photon-subtracted TMS states [41].
Figure 2 shows $\Delta\mathcal{E}_{\text{rel}}$ as a function of the squeezing and fluctuation parameters $z$ and $k$ of the initial Gaussian TMS state prior to photon subtraction. Here, G-REG ($\Delta\mathcal{E}_{G}$) is computed using Eqs. (3), (5), and (6), together with the covariance matrix derived in [59]. As a separability benchmark we employ the Shchukin–Vogel (SV) criterion [42], whose negativity suffices to certify entanglement. The maximum value of $\Delta\mathcal{E}_{\text{rel}}$ observed within the separable region (above the SV boundary) is then taken as a threshold, indicated by the dashed line in the figure.
Lemma 4.
For the non-Gaussian family of photon-subtracted two-mode squeezed states, $\Delta\mathcal{E}_{\textup{rel}}\gtrapprox 1.11$ implies entanglement.
A detailed proof is given in Appendix G.3 [49]. We emphasize that the separability bounds we establish for non-Gaussian states are inherently class dependent: a bound derived for one family cannot, in general, be used as an entanglement witness for another. It therefore remains an open problem whether a universal nontrivial bound for all non-Gaussian separable states exists. If such a bound were found, G-REG could then serve as a fully general tool for entanglement witnessing in the non-Gaussian regime, independent of prior knowledge of the underlying state family.
@@
Figure 2: Gaussian relative ergotropic gap of photon-subtracted two-mode squeezed states as a function of the fluctuation parameter $k$ of the initial thermal state and the Gaussian squeezing parameter $z$. The region below the solid line is formed by entangled states as witnessed by the Shchukin-Vogel (SV) condition, while the region below the dashed line indicates entangled states witnessed by our G-REG-based criterion [Lemma 4].
@@
Discussion.– In this Letter, we characterize correlations in continuous-variable systems through thermodynamic processes. For bipartite Gaussian states, we demonstrate that correlation guarantees a strictly positive ergotropic gap, revealing an energy-based signature of correlations independent of entropy. Physically, this means that energy can be locked into correlations, making it inaccessible via local operations alone. For pure states, this yields an entanglement measure equivalent to the quantum mutual information. However, in mixed states, the ergotropic gap grows while correlations decrease due to the unbounded energy spectrum. To address this, we introduce the relative ergotropic gap (REG), a normalized quantity that is functionally independent of the quantum mutual information and more generally of entropic quantifiers. We derive analytical bounds on REG that distinguish entangled from separable states and serve as a necessary and sufficient separability criterion for a large class of states, paralleling the PPT test. Notably, when energy measurements are available [60], this could provide a more laboratory-friendly entanglement indicator than full state tomography. Moreover, we show that REG may find broader applicability outside the Gaussian regime — an open direction for future study. Overall, our work establishes REG as an energy-based, experimentally accessible witness of continuous-variable quantum correlations, bridging thermodynamics and information-theoretic perspectives, and offering a novel tool for entanglement detection and characterization in near-term quantum optical technologies.
@@ Acknowledgements.
Acknowledgements.– This project has been funded by the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program), and European Union (Quantera Veriqtas). F.C. acknowledges funding from the European Union (EQC, 101149233). G.A. acknowledges funding from the UK Research and Innovation (UKRI EPSRC Grant No. EP/X010929/1). M.A. acknowledges funding from the European Union (QURES, 101153001). The views and opinions expressed in this Letter are, however, the sole responsibility of the author(s).
Data availability.— The data supporting the findings of this study are available in Ref. [61].
Appendix A Separability of Gaussian two-mode states
@@
$\sigma_{sf}=\begin{pmatrix}a&0&c_{1}&0\\ 0&a&0&c_{2}\\ c_{1}&0&b&0\\ 0&c_{2}&0&b\end{pmatrix},\quad\sigma_{A}=a\mathbb{I}_{2},\quad\sigma_{B}=b\mathbb{I}_{2},\quad\sigma_{AB}=\begin{pmatrix}c_{1}&0\\ 0&c_{2}\end{pmatrix}$,
@@
Appendix B Proof of Theorem 1
@@
Appendix C Parametrization of Gaussian two-mode mixed states
@@
In what follows, we will adopt the xpxpxpxp convention for mode ordering for the definition of the matrices appearing in the previous expression; that is, we will take the quadrature vector as $\hat{\vec{r}}=(\hat{x}_{A},\hat{p}_{A},\hat{x}_{B},\hat{p}_{B})^{T}$ as opposed to the alternative ordering $\hat{\vec{r}}=(\hat{x}_{A}\hat{x}_{B},\hat{p}_{A},\hat{p}_{B})^{T}$. The matrix
@@
$\omega\equiv\omega_{A}$: the natural oscillating frequency of mode AA
@@
For the example of two-mode squeezed (TMS) states, discussed in the main text to illustrate the genuinely distinct behaviour of Continuous variable (CV) and DV ones, the above parametrization yields the following form of the covariance matrix, which is itself in standard form:
$\small\sigma_{\text{TMS}}=\left[\begin{matrix}k\cosh{\left(2r\right)}\mathbb{I}&k\sinh{\left(2r\right)}\sigma_{z}\\ k\sinh{\left(2r\right)}\sigma_{z}&k\cosh{\left(2r\right)}\mathbb{I}\end{matrix}\right],\,\,\text{where}\begin{cases}k=\coth\left(\mbox{$\frac{\omega}{2T}$}\right)\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}e^{-2r}=z}\end{cases}$
Appendix D Calculation of relative ergotropic gaps
@@
D.1 Gaussian two-mode states & Proof of Lemma 1
@@
$\sigma_{\text{gp}}=G\ \sigma G^{T}\$(25)
@@
Substituting expressions 30,31 and 20 into Eq. 29, we derive the explicit covariance matrix of the local passive state, whose energy, computed through Eq.3, yields:
$E_{lp}=\frac{\omega}{2}\Bigg[\sqrt{(k+\gamma)^{2}\cos^{4}\theta+(k-\gamma)^{2}\sin^{4}\theta+(k^{2}-\gamma^{2})\cos^{2}\theta\sin^{2}\theta\bigg(\frac{z_{1}^{2}+z_{2}^{2}}{z_{1}z_{2}}\bigg)}+\alpha\sqrt{(k-\gamma)^{2}\cos^{4}\theta+(k+\gamma)^{2}\sin^{4}\theta+(k^{2}-\gamma^{2})\cos^{2}\theta\sin^{2}\theta\bigg(\frac{z_{1}^{2}+z_{2}^{2}}{z_{1}z_{2}}\bigg)}-(1+\alpha)\Bigg]$
@@
D.2 REG and Quantum Discord in the Gaussian Regime
Lemma 1 established that the ergotropic gap acts as a faithful quantifier of total correlations. However, since this quantity diverges for alomost correlation free states, it does not qualify as a well-defined correlation quantifier. By introducing the regularized form—the relative ergotropic gap (REG)—we resolve this issue and obtain a faithful and well defined total correlation quantifier. Our central motivation, however, goes beyond total correlations: we seek a thermodynamically inspired, entropy-free criterion that can not only capture classical correlations but also reveal genuinely quantum features of a state. In particular, we ask whether REG can distinguish classical from quantum correlations, thereby serving as a witness of quantumness.
Before characterizing quantumness, it is useful to recall how quantum correlation is defined via the structure of the state space. Restricting to bipartite systems, any state can be classified as follows:
  • Classical–Classical (CC): Both subsystems are diagonal in some orthonormal bases.
[…]
  • Classical–Quantum (CQ): The first subsystem is diagonal in an orthonormal basis; the second is not.
  • Quantum–Classical (QC): The second subsystem is diagonal in an orthonormal basis; the first is not.
  • Quantum–Quantum (QQ): Separable states that are neither CC, CQ, nor QC.
  • Entangled (ENT): States that are not separable.
A state in the CC class is entirely classical, as both subsystems consist of distinguishable states. This observation underlies the concept of quantum discord [65, 66, 67], which captures quantumness beyond entanglement by measuring the difference in mutual information across local measurements. Quantum discord is inherently asymmetric: for example, if Alice’s subsystem lacks a classical description (i.e., is not diagonal in any orthonormal basis), then the discord from Alice to Bob ($D_{A}$) is non-zero.
The discord for the five categories of states is summarized below:
[…]
Research in DV systems has explored the relationship between the ergotropic gap and discord [21, 20, 31]. Specifically, [21] shows that for a non-degenerate Hamiltonian, the ergotropic gap $\Delta\mathcal{E}(\rho)$ is strictly positive for all non-CC states:
$\neg(\text{CC})\Rightarrow\Delta\mathcal{E}(\rho)>0.$
This leads to the following correspondence:
[…]
Note that even product states in the CC category may exhibit a non-zero ergotropic gap in DV systems. However, this does not hold when the Hamiltonian is degenerate [22].
We now extend this analysis to continuous-variable (CV) systems with REG as quantifier, focusing on two-mode Gaussian states. According to Adesso and Datta [68], such states fall into only two categories:
[…]
Combining this classification with our Lemma 1, we obtain the following relationship for Gaussian states:
[…]
In the Gaussian regime, the absence of classical correlations implies that all non-product states exhibit quantum correlations. Thus, $\Delta\mathcal{E}_{\text{rel}}(\rho)=0$ only when the state is a product. Unlike DV systems, Gaussian product states cannot yield a non-zero ergotropic gap—a limitation imposed by the Gaussianity constraint. Non-Gaussian states, however, may exhibit such phenomena [46].
In summary, there is a robust connection between quantum discord and the ergotropic gap in Gaussian systems. While mutual information, discord, and the relative ergotropic gap each serve as valid indicators of quantum correlations, the ergotropic gap uniquely provides an energy-based characterization that complements traditional entropy-based measures.
Appendix E Proof of Theorem 2 and Corollary 1
@@
Part (a). First, we note that by expansion of the binomial squares $(k+\gamma)^{2}$ and $(k-\gamma)^{2}$, the sum of square roots in the expression of the relative ergotropic gap (Eq. 33) yields the following structure
@@
Figure 3 illustrates the meaning of these two bounds. Relative ergotropic gap is shown for $N=5\times 10^{2}$ randomly generated two-mode Gaussian states. The blue dashed line represents the minimum $\Delta\mathcal{E}_{\text{rel}}$ for entangled states, while the upper bound for separable states is displayed in dashed red. Since the two bounds $\mathcal{B}^{\text{ent}}_{\mathrm{min}}$ and $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ are dependent on $k,\gamma$ and $\alpha$, the values of these parameters have been fixed so a single threshold holds for all of the simulated states. Neither entanglement nor separability of states with REG lying in between the two lines can be witnessed by means of our criterion. However, any value of REG above the red line or below the blue one can unambiguously certify that the state is entangled or separable, respectively.
@@
If $\rho_{G}$ has equal fluctuation factors $k_{A},k_{B}$ for both modes, then $k=k_{A}=k_{B}$ and $\gamma=0$. In that case, $\mathcal{B}^{\text{sep}}_{\mathrm{max}}=\mathcal{B}^{\text{ent}}_{\mathrm{min}}$. From part (a) of Theorem 2, we derive that Eq.9 is necessary for separability. Sufficiency follows from part (b). ∎
Appendix F Independency of relative ergotropic gap
F.1 Proof of Lemma 2